# MGE-LDM: Joint Latent Diffusion for Simultaneous Music Generation and Source Extraction

### v1

@@ 5.2 Source Extraction
We evaluate text-queried source extraction using the Log Mel L1 distance, following MSG-LD [35], due to the inherent phase mismatch in latent-domain models that complicates waveform-domain evaluation. Table 3 presents results for a variety of stems across the $\textbf{S}_{A}\,$, $\textbf{S}_{B}\,$, $\textbf{M}_{u}\,$, and $\textbf{M}_{o}\,$test sets. Alongside MSDM and MSG-LD, we compare with two additional baselines: Demucs [41], a strong waveform-domain separation model trained on fixed stems; and AudioSep [52], a recent system that enables natural language-driven extraction. For fixed-stem baselines (e.g., MSDM, MSG-LD, Demucs), we report both in-distribution results and out-of-distribution generalization where possible (e.g., bass in $\textbf{M}_{u}\,$).
Our model $\mathcal{T}_{1}\,$, trained solely on $\textbf{S}_{A}\,$, performs strongly on the canonical Slakh stems (bass, drums, guitar, piano), outperforming MSG-LD on all but bass. However, it generalizes poorly to less common stems and real-world recordings. By expanding the training set to encompass the full Slakh2100 dataset, model $\mathcal{T}_{2}\,$achieves improved performance on categories such as chromatic percussion, organ, synth lead, and synth pad, demonstrating the importance of broader intra-domain coverage.
@@
Overall, MGE-LDM delivers high-fidelity, class-agnostic extraction with strong performance across a wide range of instrument types and recording domains, highlighting its effectiveness for text-driven music source extraction in both synthetic and real-world settings.
@@ Appendix A Background
$\boldsymbol{v}_{\tau}=\frac{d\mathbf{z}_{\tau}}{d\phi_{\tau}}=\frac{d\cos(\phi_{\tau})}{d\phi_{\tau}}\mathbf{z_{0}}+\frac{d\sin(\phi_{\tau})}{d\phi_{\tau}}\boldsymbol{\epsilon}=-\sin(\phi_{\tau})\textbf{z}_{0}+\cos(\phi_{\tau})\boldsymbol{\epsilon}=\alpha_{\tau}\boldsymbol{\epsilon}-\beta_{\tau}\mathbf{z}_{0},$ (15)
@@
Recovering $\textbf{z}_{0}$ and $\boldsymbol{\epsilon}$. Using the definition of $\boldsymbol{v}_{\tau}$, we can rearrange the terms to recover the clean latent by:
@@ A.2 Canonical Inpainting in Score-Based Models
The core idea behind inpainting in score-based generative models is to estimate the score of the unknown region conditioned on known region [11, 33].
@@
[…]
@@
where $[\textbf{z}_{\tau}^{\Gamma};\hat{\textbf{z}}_{\tau}^{\Omega}]$ denotes a composite latent vector such that the known region is replaced by $\hat{\textbf{z}}_{\tau}^{\Omega}$ while the unknown region remains as $\textbf{z}_{\tau}^{\Gamma}$, adopting the same notation as Song et al. [11].
@@ A.3 RePaint
Lugmayr et al. proposed RePaint [83], a resampling-based mechanism that improves score-based inpainting by repeating the diffusion process across multiple forward–reverse cycles. Their key insight is that, in conventional inpainting (as described in Equation (20)), the sampled noise for the known region is independent of the generated (inpainted) region. This lack of synchronization can lead to semantic inconsistencies and disharmony between known and unknown parts of the sample.
To address this, RePaint introduces a resampling mechanism during generation. At each denoising timestep, the algorithm alternates between one reverse diffusion step and one forward diffusion step, repeating this cycle $U$ times. These micro-steps refine the sampling distribution and can have the effect of partially marginalizing over the known region at noise level $\tau$ in Equation (20), thereby reducing the approximation error inherent in conditional score estimation. This iterative resampling procedure improves consistency but incurs higher computational cost, as each denoising step requires multiple forward–reverse passes–making RePaint significantly more expensive than standard inpainting methods.
@@
8:$\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
9:$\boldsymbol{\tau}_{T}\leftarrow\tau_{T}(1-m)$
10:$\textbf{z}_{\boldsymbol{\tau}_{T}}=m\textbf{z}_{0}^{\text{known}}+(1-m)\boldsymbol{\epsilon}$
@@ Appendix B Adaptive Timestep Conditioning and Score Approximation
Conventional inpainting approaches approximate the conditional score in Eq. (18) using a single sampled estimate of the known latents. This corresponds to a high-variance Monte Carlo estimate of the expectation over $q_{\tau}(\textbf{z}_{\tau}^{\Omega},|\textbf{z}_{0}^{\Omega})$, which may lead to instability–especially at high noise levels.
By contrast, the adaptive timestep model proposed in Section 3.4–inspired by TD-Paint [86]–circumvents this marginalization by training the model to directly approximate the conditional score.
@@
[…]
@@
as in Section 3.4. The model is trained using the velocity objective in Eq. (12), restricted to the unknown region. This allows the denoiser to explicitly learn the conditional score on $\textbf{z}_{\tau}^{\Gamma}$, avoiding the need for the stochastic marginalization and improving accuracy in conditional inpainting tasks. The full sampling procedure is detailed in Algorithm 2.
@@ Appendix C Iterative Generation
Let =$\mathcal{I}=\{c^{(s)}_{i}\}_{i}$ be an (ordered) set of CLAP embeddings corresponding to the desired instrument description. At the first iteration ($i=1$), we generate an initial source latent $\hat{z}^{(s)}_{1}$ by sampling with the model conditioned only on the prompt $c^{(s)}_{1}$:
@@
and set $z^{(u)}_{1}=\hat{z}^{(s)}_{1}$ as t he initial submixture latent. For each subsequent iteration $i=2,...,|\mathcal{I}|$, we follow the iterative imputation strategy of partial generation, treating the current submixture as the accumulated sum of decoded sources from the previous steps.
Finally, the full mixture waveform is constructed via decoding and summing the generated source latents:
@@ D.1 Datasets
Slakh2100 is derived from the Lakh MIDI Dataset v0.1 [93] and contains synthesized tracks rendered with sample-based virtual instruments. It comprises 2100 songs divided into training (1500), validation (375), and test (225) splits, totaling approximately 145 hours of audio. It includes a wide variety of instrument classes (e.g., bass, drums, guitar, piano, strings, synth pad, etc.). We adopt the naming $\textbf{S}_{A}\,$to denote a subset containing only bass, drums, guitar, and piano–the four classes used by MSDM and MSG-LD–and $\textbf{S}_{B}\,$to denote the complementary subset of remaining stems. We follow the official dataset splits provided by Slakh2100 for training, validation, and testing.
MUSDB18 consists of 150 real-world music recordings with four stems: drums, bass, other, and vocals. We use all 100 tracks from the official training split for training, and the 50-track test split for evaluation. The total dataset length is approximately 10 hours.
MoisesDB comprises 240 songs (14 hours total) contributed by 47 artists across 12 genres. Each stem in the song is annotated with a two-tier stem taxonomy. Each track is decomposed into its constituent sources and annotated using a two-level hierarchical taxonomy of stem classes. We aggregate all second-level tracks into their corresponding top-level class. Among the 11 stem classes, we evaluate only the 7 unambiguous stems (e.g., bass, percussion, vocals, etc.). For evaluation, we randomly sample 24 tracks (10%) as the test set and use the remaining tracks for training.
@@ D.2 Baseline Implementations
All baseline metrics are recomputed on our test splits for fair comparison. The following models are used:
@@ D.3 Model Architecture and Training
Autoencoder. We adopt the VAE-based architecture from Stable Audio [71], with a downsampling ratio of 2048, yielding a 7.8125 Hz latent resolution and 64 latent channels. We train the autoencoder on all training subsets from Slakh2100, MUSDB18, and MoisesDB using 16 kHz mono audio for 600K steps with batch size 16.
Diffusion Model. In practice, the three latent representations are concatenated along the channel dimension, such that the input to the diffusion model becomes $\textit{Concat}[z^{(m)},z^{(u)},z^{(s)}]\in\mathbb{R}^{3C\times L}$. We use a DiT backbone [80] with 24 transformer blocks with 48 heads, and a projected latent dimension of 1536 (3 tracks × 512 each). Timestep embeddings are prepended to the input vector of the transformer. CLAP embeddings for each track are processed by independent projection layers (without weight sharing) to produce scale and shift parameters for AdaIN-style conditioning [94]. These are applied group-wise via GroupNorm within each DiT layer to modulate the corresponding track-specific activations. Text embeddings are obtained from CLAP [55] using the checkpoint music_audioset_epoch_15_esc_90.14.pt via the laion-clap library. [^6] https://github.com/LAION-AI/CLAP Our implementation builds upon the official stable-audio-tools repository from Stability AI [^7] https://github.com/Stability-AI/stable-audio-tools and the training framework from friendly-stable-audio-tools. [^8] https://github.com/yukara-ikemiya/friendly-stable-audio-tools All models were trained on a single NVIDIA RTX 6000 GPU (48 GB memory).
@@ Appendix E Ablation Study
This section presents ablation experiments designed to further analyze key components of our framework. Unless otherwise specified, all models are trained on the combined $\textbf{S}_{A}\,$++\ +$\textbf{S}_{B}\,$dataset. Each of our models is evaluated with the same configuration as in Section 5, using $T=250$ denoising steps during sampling.
Table 5: Source extraction performance of RePaint-based methods vs. MGE-LDM. Metrics are reported as Log-Mel L1 distance $\downarrow$. $T$ indicates the number of reverse timesteps, and $U$ specifies the number of denoising operations per reverse step (i.e., $U{-}1$ intermediate resampling steps). The case $U=1$ corresponds to the canonical single-sample conditional score approximation [11], as described in Appendix A.2. MGE-LDM uses adaptive timestep conditioning without resampling.
@@ E.1 Comparison with Canonical Inpainting Methods
We assess the effectiveness of our adaptive timestep conditioning strategy by comparing it against two prior approaches: the canonical one-sample conditional score approximation (Appendix A.2) and the RePaint method [83] (Appendix A.3). Table 5 reports the results for the source extraction task.
In RePaint, $U$ denotes the number of denoising steps performed per reverse timestep: one denoising step followed by $U{-}1$ forward (resampling) steps. As a result, the total number of denoising steps becomes $T\times U$ during the full inpainting process. Note that setting $U=1$ recovers the canonical single-sample estimator in Eq. (20).
@@
Interestingly, we also observe that RePaint configurations with larger total denoising steps – such as $T=250,U=2$ and $T=250,U=4$ – consistently underperform compared to $T=25,U=10$, across all stems in both $\textbf{S}_{A}\,$and $\textbf{S}_{B}\,$. This suggests that, for inpainting tasks, accurately modeling the conditional score at each timestep is more critical than simply increasing the number of reverse steps. As RePaint approximates the conditional score by marginalizing over perturbed conditions via resampling, performance benefits are observed primarily through increased resampling ($U$), not longer trajectories ($T$).
@@
[…]
Table 6: Total generation performance across modeling variants. Metrics are reported as FAD $\downarrow$. All models are trained on $\textbf{S}_{A}\,$++\,+$\textbf{S}_{B}\,$. The baseline model uses uniform (non-adaptive) timesteps across all tracks. MGE variants apply adaptive timestep conditioning and test the impact of normalizations and CFG dropout rates. Values in parentheses indicate generation conditioned on the text prompt "The sound of the bass, drums, guitar, and piano".
@@
- CFG dropout $p{=}0.5$ 3.12 (2.67) 0.58 5.43 4.82
As shown in Table 6, comaprining non-adaptive uniform timstep basline model and our model, both models achieve comparable FAD scores, indicating that timestep adaptation preserves generation performance under uniform scheduling while providing significant advantages for inpainting tasks.

### v2

@@ 5.2 Source Extraction
We evaluate text-queried source extraction using the Log Mel L1 distance, following MSG-LD [36], due to the inherent phase mismatch in latent-domain models that complicates waveform-domain evaluation. Table 3 presents results for a variety of stems across the $\textbf{S}_{A}$, $\textbf{S}_{B}$, $\textbf{M}_{u}$, and $\textbf{M}_{o}$ test sets. Alongside MSDM and MSG-LD, we compare with two additional baselines: HDemucs [20], a strong waveform- and spectrogram-domain separation model trained on fixed stems; and AudioSep [53], a recent system that enables natural language-driven extraction. For fixed-stem baselines (e.g., MSDM, MSG-LD, HDemucs), we report both in-distribution results and out-of-distribution generalization where possible (e.g., bass in $\textbf{M}_{u}$).
Our model $\mathcal{T}_{1}$, trained solely on $\textbf{S}_{A}$, performs strongly on the canonical Slakh stems (bass, drums, guitar, piano), outperforming MSG-LD on all but bass. However, it generalizes poorly to less common stems and real-world recordings. By expanding the training set to encompass the full Slakh2100 dataset, model $\mathcal{T}_{2}$ achieves improved performance in categories such as chromatic percussion, organ, synth lead, and synth pad, demonstrating the importance of broader intra-domain coverage.
@@
Overall, MGE-LDM delivers robust, class-agnostic extraction with strong performance across a wide range of instrument types and recording domains, highlighting its effectiveness for text-driven music source extraction in both synthetic and real-world settings.
6 Limitations
While MGE-LDM provides a flexible, class-agnostic framework for multi-track audio modeling, several limitations remain. First, all experiments are conducted using 16 kHz monaural audio, which constrains upper-frequency resolution and omits spatial cues, thereby limiting realism for high-fidelity or stereo music applications. Second, the model relies on CLAP-based semantic conditioning, which introduces a modality gap between text and audio [99]. This can occasionally lead to semantic drift during extraction
Third, although MGE-LDM reduces dependence on fixed instrument classes, it still requires multi-stem supervision during training. This dependency restricts applicability to fully unlabeled or large-scale web audio collections. Fourth, training on MUSDB18 alone with the same number of iterations as other configurations leads to overfitting, likely due to the limited duration (approximately 10 hours) of its training split. This highlights the challenge of achieving robust performance in low-resource multi-track settings.
Finally, our model is trained using triplets $(\textit{mix},\textit{submix},\textit{source})$ that satisfy $\textit{mix}=\textit{submix}+\textit{source}$ in waveform space; however, the latent diffusion process does not enforce an explicit additivity constraint for generated triplets. We believe this omission contributes directly to hallucination penomena, where the model extracts source absent from the mixture. Postolache et al. [27] addresses a related issue by enforcing additivity in a discrete VQ-VAQE latent space, estimating the joint likelihood of two sources by counting codebook co-occurrences — effectively modeling $p(z_{\text{mix}}|z_{\text{src}_{1}},z_{\text{src}_{2}})$, where $z_{*}$ are quantized latent codes. Our current pipeline, however, operates in a continuous latent space, which precludes the direct use of such discrete bin-counting methods. Adapting this latent-domain likelihood formulation to continuous spaces, for example, by designing suitable regularizers or adopting a VQ-VAE-based encoder with discrete diffusion [100, 101, 102] or MaskGiT [65]-style generation, represents a promising direction for future work.
@@ 7 Conclusion
Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) [No. RS-2024-00461617, 90%], Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University), 10%] Additionally, the GPUs were partly supported by the National IT Industry Promotion Agency (NIPA)’s high-performance computing support program in 2025.
@@ Appendix A Background
$\boldsymbol{v}_{\tau}=\frac{\partial\mathbf{z}_{\tau}}{\partial\phi_{\tau}}=\frac{\partial\cos(\phi_{\tau})}{\partial\phi_{\tau}}\mathbf{z_{0}}+\frac{\partial\sin(\phi_{\tau})}{\partial\phi_{\tau}}\boldsymbol{\epsilon}=-\sin(\phi_{\tau})\textbf{z}_{0}+\cos(\phi_{\tau})\boldsymbol{\epsilon}=\alpha_{\tau}\boldsymbol{\epsilon}-\beta_{\tau}\mathbf{z}_{0},$ (15)
@@
Recovering $\textbf{z}_{0}$ and $\boldsymbol{\epsilon}$. Using the definition of $\boldsymbol{v}_{\tau}$ and recalling the perturbation process defined in Eq. (2),
$\textbf{z}_{\tau}=\alpha_{\tau}\textbf{z}_{0}+\beta_{\tau}\boldsymbol{\epsilon}\,$ (16)
we can rearrange the terms to recover the clean latent by:
@@ A.2 Canonical Inpainting in Score-Based Models
The core idea behind inpainting in score-based generative models is to estimate the score of the unknown region conditioned on the known region [11, 34].
@@
$q_{\tau}(\textbf{z}_{\tau}^{\Gamma}|\textbf{z}_{0}^{\Omega})$ $=\int q_{\tau}(\textbf{z}_{\tau}^{\Gamma},\textbf{z}_{\tau}^{\Omega}|\textbf{z}_{0}^{\Omega})d\textbf{z}_{\tau}^{\Omega}$
$=\int q_{\tau}(\textbf{z}_{\tau}^{\Gamma}|\textbf{z}_{\tau}^{\Omega},\textbf{z}_{0}^{\Omega})q_{\tau}(\textbf{z}_{\tau}^{\Omega}|\textbf{z}_{0}^{\Omega})d\textbf{z}_{\tau}^{\Omega}$
@@
where $\textbf{z}_{\tau}^{\Gamma};\hat{\textbf{z}}_{\tau}^{\Omega}]$ denotes a composite latent vector, such that the known region is replaced by $\hat{\textbf{z}}_{\tau}^{\Omega}$, while the unknown region remains as $\textbf{z}_{\tau}^{\Gamma}$, adopting the same notation as Song et al. [11].
@@ A.3 RePaint
Lugmayr et al. proposed RePaint [89], a resampling-based mechanism that improves score-based inpainting by repeating the diffusion process across multiple forward–reverse cycles. Their key insight is that, in conventional inpainting (as described in Equation (21)), the sampled noise for the known region is independent of the generated (inpainted) region. This lack of synchronization can lead to semantic inconsistencies and disharmony between the known and unknown parts of the sample.
To address this, RePaint introduces a resampling mechanism during generation. At each denoising timestep, the algorithm alternates between one reverse diffusion step and one forward diffusion step, repeating this cycle $U$ times. These micro-steps refine the sampling distribution and can have the effect of partially marginalizing over the known region at noise level $\tau$ in Equation (21), thereby reducing the approximation error inherent in conditional score estimation. This iterative resampling procedure improves consistency but incurs a higher computational cost, as each denoising step requires multiple forward–reverse passes, making RePaint significantly more expensive than standard inpainting methods.
@@
[…]
@@ Appendix B Adaptive Timestep Conditioning and Score Approximation
Conventional inpainting approaches approximate the conditional score in Eq. (19) using a single sampled estimate of the known latents. This corresponds to a high-variance Monte Carlo estimate of the expectation over $q_{\tau}(\textbf{z}_{\tau}^{\Omega}|\textbf{z}_{0}^{\Omega})$, which may lead to instability — especially at high noise levels.
By contrast, the adaptive timestep model proposed in Section 3.4, inspired by TD-Paint [92], circumvents this marginalization by training the model to directly approximate the conditional score.
@@
$q_{\tau}(\textbf{z}_{\tau}^{\Gamma}|\textbf{z}_{0}^{\Omega})$ $=\int q_{\tau}(\textbf{z}_{\tau}^{\Gamma},\mathbf{x}_{0}^{\Gamma}|\textbf{z}_{0}^{\Omega})d\mathbf{x}_{0}^{\Gamma}$
$=\int q_{\tau}(\textbf{z}_{\tau}^{\Gamma}|\mathbf{x}_{0}^{\Gamma},\textbf{z}_{0}^{\Omega})q(\mathbf{x}_{0}^{\Gamma}|\textbf{z}_{0}^{\Omega})d\mathbf{x}_{0}^{\Gamma}$
@@
as in Section 3.4. The model is trained using the velocity objective in Eq. (12), restricted to the unknown region. This allows the denoiser to explicitly learn the conditional score on $\textbf{z}_{\tau}^{\Gamma}$, avoiding the need for stochastic marginalization and improving accuracy in conditional inpainting tasks. The full sampling procedure is detailed in Algorithm 2.
@@ Appendix C Iterative Generation
Let $\mathcal{I}=\{c^{(s)}_{i}\}_{i}$ be an (ordered) set of CLAP embeddings corresponding to the desired instrument description. At the first iteration ($i=1$), we generate an initial source latent $\hat{z}^{(s)}_{1}$ by sampling with the model conditioned only on the prompt $c^{(s)}_{1}$:
@@
[…]
Finally, the full mixture waveform is constructed by decoding and summing the generated source latents:
@@ D.1 Datasets
Slakh2100 is derived from the Lakh MIDI Dataset v0.1 [103] and contains synthesized tracks rendered with sample-based virtual instruments. It comprises 2100 songs divided into training (1500), validation (375), and test (225) splits, totaling approximately 145 hours of audio. It includes a wide variety of instrument classes (e.g., bass, drums, guitar, piano, strings, synth pad, etc.). We adopt the naming $\textbf{S}_{A}$ to denote a subset containing only bass, drums, guitar, and piano, the four classes used by MSDM and MSG-LD, and $\textbf{S}_{B}$ to denote the complementary subset of the remaining stems. We follow the official dataset splits provided by Slakh2100 for training, validation, and testing.
MUSDB18 consists of 150 real-world music recordings with four stems: drums, bass, other, and vocals. We use all 100 tracks from the official training split for training and the 50-track test split for evaluation. The total dataset length is approximately 10 hours.
MoisesDB comprises 240 songs (14 hours total) contributed by 47 artists across 12 genres. Each stem in the song is annotated with a two-tier stem taxonomy. Each track is decomposed into its constituent sources and annotated using a two-level hierarchical taxonomy of stem classes. We aggregate all second-level tracks into their corresponding top-level classes. Among the 11 stem classes, we evaluate only the 7 unambiguous stems (e.g., bass, percussion, vocals, etc.). For evaluation, we randomly sample 24 tracks (10%) as the test set and use the remaining tracks for training.
@@ D.2 Baseline Implementations
All baseline metrics are recomputed on our test splits for a fair comparison. The following models are used:
@@ D.3 Model Architecture and Training
Autoencoder. We adopt the VAE-based architecture from Stable Audio [72], with a downsampling ratio of 2048, yielding a 7.8125 Hz latent resolution and 64 latent channels. We train the autoencoder on all training subsets from Slakh2100, MUSDB18, and MoisesDB using 16 kHz mono audio for 600K steps with a batch size of 16.
Diffusion Model. In practice, the three latent representations are concatenated along the channel dimension, such that the input to the diffusion model becomes $\textit{Concat}[z^{(m)},z^{(u)},z^{(s)}]\in\mathbb{R}^{3C\times L}$. We use a DiT backbone [86] with 24 transformer blocks with 48 heads, and a projected latent dimension of 1536 (3 tracks × 512 each). Timestep embeddings are prepended to the input vector of the transformer. CLAP embeddings for each track are processed by independent projection layers (without weight sharing) to produce scale and shift parameters for AdaIN-style conditioning [104]. These are applied group-wise via GroupNorm within each DiT layer to modulate the corresponding track-specific activations. Text embeddings are obtained from CLAP [56] using the checkpoint music_audioset_epoch_15_esc_90.14.pt via the laion-clap library. [^5] https://github.com/LAION-AI/CLAP Our implementation builds upon the official stable-audio-tools repository from Stability AI [^6] https://github.com/Stability-AI/stable-audio-tools and the training framework from friendly-stable-audio-tools. [^7] https://github.com/yukara-ikemiya/friendly-stable-audio-tools All models were trained on a single NVIDIA RTX 6000 GPU (48 GB of memory).
@@ Appendix E Ablation Study
This section presents ablation experiments designed to further analyze the key components of our framework. Unless otherwise specified, all models are trained on the combined $\textbf{S}_{A}$$\textbf{S}_{B}$ dataset. Each of our models is evaluated with the same configuration as in Section 5, using $T=250$ denoising steps during sampling.
Table 5: Source extraction performance of RePaint-based methods vs. MGE-LDM. Metrics are reported as Log-Mel L1 distance $\downarrow$. $T$ indicates the number of reverse timesteps, and $U$ specifies the number of denoising operations per reverse step (i.e., U1U{-}1 intermediate resampling steps). The case $U=1$ corresponds to the canonical single-sample conditional score approximation [11], as described in Appendix A.2. MGE-LDM uses adaptive timestep conditioning without resampling.
@@ E.1 Comparison with Canonical Inpainting Methods
We assess the effectiveness of our adaptive timestep conditioning strategy by comparing it against two prior approaches: the canonical one-sample conditional score approximation (Appendix A.2) and the RePaint method [89] (Appendix A.3). Table 5 reports the results of the source extraction task.
In RePaint, $U$ denotes the number of denoising steps performed per reverse timestep: one denoising step followed by U1U{-}1 forward (resampling) steps. As a result, the total number of denoising steps becomes $T\times U$ during the full inpainting process. Note that setting $U=1$ recovers the canonical single-sample estimator in Eq. (21).
@@
Interestingly, we also observe that RePaint configurations with larger total denoising steps — such as $T=250,U=2$ and $T=250,U=4$ — consistently underperform compared to $T=25,U=10$, across all stems in both $\textbf{S}_{A}$ and $\textbf{S}_{B}$. This suggests that, for inpainting tasks, accurately modeling the conditional score at each timestep is more critical than simply increasing the number of reverse steps. As RePaint approximates the conditional score by marginalizing over perturbed conditions via resampling, performance benefits are observed primarily through increased resampling ($U$), not longer trajectories ($T$).
@@
A potential concern is whether optimizing for adaptive timestep-conditional inference might degrade generation quality when using uniform timestep schedules across tracks. To assess this, we evaluate our adaptive timestep model with a uniform timestep vector $\boldsymbol{\tau}=(\tau,\tau,\tau)$, which corresponds to the total generation task, and compare it to a baseline trained with non-adaptive, shared timesteps.
Table 6: Total generation performance across modeling variants. Metrics are reported as FAD $\downarrow$. All models are trained on $\textbf{S}_{A}$$\textbf{S}_{B}$. The baseline model uses uniform (non-adaptive) timesteps across all tracks. MGE variants apply adaptive timestep conditioning and test the impact of normalizations and CFG dropout rates. Values in parentheses indicate generation conditioned on the text prompt "The sound of the bass, drums, guitar, and piano".
@@
- CFG dropout p=0.5p{=}0.5 3.12 (2.67) 0.58 5.43 4.82
As shown in Table 6, comparing the non-adaptive uniform timestep baseline model and our model, both models achieve comparable FAD scores, indicating that timestep adaptation preserves generation performance under uniform scheduling while providing significant advantages for inpainting tasks.